OpenStudy (anonymous):

A tent is shaped like a pyramid. It has a square base measuring 8ft on a side and a height of 6ft. What is the volume of the tent?

OpenStudy (anonymous):

V = 1/3 base area * height = 64*2 = 128 cubic feet.